OpenStudy (anonymous):

solve 5/4x+1/8x=9/8+x

OpenStudy (anonymous):

5/4x+1/8x=9/8+x 11/8x = 9/8+x (simplifying LHS) 3/8x = 9/8 (subtracting x from both sides , in this case 8/8x) x = 9/8 divided by 3/8 = 3 x=3

OpenStudy (angela210793):

@gian don't u have 12/8x?

OpenStudy (angela210793):

oops no

OpenStudy (anonymous):

just convert all fractions to eighths

OpenStudy (anonymous):

so would the ans. be 3 or 4

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Plug three into the equation to see.

OpenStudy (angela210793):

yea i know...is just tht i added 2 instead of 1 :P